40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 12 17:37:11 2011, Jenne, Update, SUS, ETMY hopefully good again 
    Reply  Fri Aug 12 20:28:13 2011, Jenne, Update, SUS, ETMY hopefully good again 
       Reply  Sat Aug 13 02:31:42 2011, kiwamu, Update, SUS, Re: ETMY hopefully good again 
          Reply  Mon Aug 15 13:16:50 2011, Jenne, Update, SUS, Re: ETMY hopefully good again 
Message ID: 5216     Entry time: Fri Aug 12 20:28:13 2011     In reply to: 5215     Reply to this: 5221
Author: Jenne 
Type: Update 
Category: SUS 
Subject: ETMY hopefully good again 

Quote:

[Jamie, Jenne]

We went in to have a look-see at ETMY since it looked stuck-ish.  Jamie noticed that the side magnet was pretty close to the teflon plates of the OSEM.  We rotated it a bit, and now its all better.  We also adjusted the OSEMs until their mid-ranges were happy.  The U's were a little low, and the L's were a little high, as if the optic were a bit pitched backward.  Anyhow, we checked that the table is level, and tweaked the OSEMs.  We're starting the free-swinging test now...

Excited all optics

Fri Aug 12 17:38:53 PDT 2011
997231148

 Hmmm.  I'm no longer convinced that ETMY is healthy.  I think that when I gave it a kick, it's bouncing against something.  I can't fit the peaks to get the input matrix.  I guess step 1 is to try giving it a smaller kick for the free swinging spectra.  But if the owl shift folk feel like it, they might have a look-see.

ELOG V3.1.3-